An artist has been commissioned to make a stained glass window in the shape of a regular octagon. The octagon must fit inside an 18-in. square space. Determine the length of each side of the octagon. Round to the nearest hundredth of an inch
Answer by fractalier(6550)

Let the side of the octagon be x.
If you draw the diagram of it, you will notice that
x + x/sqrt(2) + x/sqrt(2) = 18
since the sloping sides are hypotenuses of 45-45-90 right triangles.
Factor out x and get
x(1 + 2/sqrt(2)) = 18
x(1 + sqrt(2)) = 18
x = 18/(1+sqrt(2)) = 7.457 in